OpenStudy (anonymous):

The amount of water a 1hp electric motor can pump from a well 10m deep is (in kg/s)?

OpenStudy (anonymous):

@akash123 @ghazi @ajprincess @CliffSedge

OpenStudy (ghazi):

1hp= 746 watt=746 j/s now your question is incomplete , for how much time will the pump work? if it keeps working it will keep pumping and practically in 1s it won't do anything lol

OpenStudy (ghazi):

P.E=m*10*10=mgh= 746, you can find m ideally

OpenStudy (anonymous):

Power = work ÷ time, work = force × distance. In this case, the force is the weight of the water, and the distance is the 10m height. \(\large \rightarrow power = \frac{weight × height}{time}, P=\frac{mgh}{t}\) \(\large \rightarrow \frac{m}{t} = \frac{P}{gh}\)